using bond energies calculate the of formation of dimethyl ether from its element from the reaction below (c(s) to c(g) = 720kj) 2c 9g)+3H2 (g)+1/2 02 (g) H3C-0- CH3(g)
1
Expert's answer
2019-11-22T04:34:52-0500

"2C(g)+3H_2(g)+1/2O_2(g) \to CH_3-O-CH_3""(g)"


As 1 mole of dimethyl ether is formed; "\Delta H_{rxn}=\Delta H_f"


"\Delta H_f=\Delta H_{f(products)}-\Delta H_{f(reactants)}"

"\Delta H_{f(products)}=6*(C-H bond energy)+" "2*(C-O bond energy)"

"=6*413+2*358"

"=3194kJ"

"\Delta H_{f(reactants)}=2*\Delta H_{f(Carbon)}+0+0" (as oxygen and hydrogen are in elemental states)

"=2*720"

"=1440kJ"


"\Delta H_f=3194-1440"

"=1754kJ" (Answer)